OpenStudy (anonymous):

Find x - 3y if x = 3 and y = -2.

OpenStudy (shaik0124):

x-3y=3-3*(-2) =3+6 =9

OpenStudy (shaik0124):

@mathi

OpenStudy (shaik0124):

@mathsucks45